# Mass Translation and Reflectivity Device Solar Sail Momentum Management Using Predictive Control

### v2

@@ 1 Introduction
The concept of a solar sail introduces the opportunity for fuel-efficient, and potentially even fuel-free, inter-planetary travel and deep-space exploration [1, 2, 3]. The momentum transferred to a solar sail by solar radiation pressure (SRP) provides a promising source of propulsive thrust that can be controlled by adjusting the attitude of the solar sail, and thus the direction and magnitude of thrust. A larger area-to-mass ratio increases a solar sail’s efficiency, which motivates the desire of developing a large-scale solar sail system, such as NASA’s Solar Cruiser, which involves a sail area greater than $1600$ m2 [4, 5]. However, larger-scale lightweight solar sails are inherently flexible, resulting in undesirable structural deformations that shift the solar sail’s center of pressure (CP) and create disturbance torques that need to be rejected by an attitude control system [6, 7, 8, 9, 10]. Reaction wheels (RWs) are typically used for this purpose, but eventually suffer from an accumulation of stored angular momentum, necessitating momentum management. Standard momentum management strategies for spacecraft, such as the use of thrusters or magnetic actuation, are not well-suited for solar sails, since they either require onboard fuel or the presence of an external magnetic field. To truly take advantage of the potential of solar sails for fuel-free deep-space exploration, more creative momentum management solutions are required.
@@
To solve this challenge, Solar Cruiser uses an active mass translator (AMT) and reflectivity control devices (RCDs) as actuators to respectively generate pitch/yaw and roll torques for momentum management [11, 12]. The AMT serves as a mechanism between two portions of the spacecraft bus, allowing them to translate relative to each other in a plane. This results in a controllable shift of the solar sail’s center of mass (CM), allowing for the generation of SRP-induced torques or the “trimming” out of disturbance torques in the pitch and yaw axes. Unlike rotative tip-devices [13, 14, 15], the RCDs are stationary thin film membranes located towards the extremities of the solar sail membrane and are canted at an angle to the plane of the sail membrane. They generate roll torques about the axis normal to the sail by adjusting their reflectivity when a voltage is applied to them. A challenge in the operation of RCDs is that they operate in an on-off fashion, either generating no torque or a constant magnitude torque in the positive or negative roll direction. Solar Cruiser’s attitude determination and control system includes control subsystems for momentum management that command the AMT and RCDs [16]. The current design of this momentum management system involves decoupled proportional-integral-derivative (PID) controllers to actuate the two axes of the AMT based on the momentum stored in the pitch/yaw axis RWs, as well as an on-off actuation strategy for the RCDs [11, 16]. Although this simple control strategy was shown to successfully perform momentum management, it neglects the dynamic coupling between the AMT and RCD actuators, which can result in undesirable performance bordering on closed-loop instability [11]. Additionally, the on-off nature of the RCDs is handled in the approach from [11, 16] using activation/deactivation thresholds, which creates additional nonlinearities in the feedback system, affecting closed-loop performance and stability properties. There is a pressing need for a momentum management control policy that is capable of explicitly accounting for the nonlinear, coupled dynamics involved in the AMT and RCD actuators, as well as the practical constraints associated with their operation.
@@
In this work, MPC is proposed for solar sail momentum management as a means to account for the environmental disturbance torques acting on the solar sail, as well as the coupled nature of the actuators and the nonlinearities of the system’s dynamics. The unique actuation properties of the AMT and RCDs make this MPC formulation challenging, resulting in an optimization problem that is naturally non-convex and unsuitable for implementation on flight hardware. To remedy this issue, we first derive a dynamic model of an AMT equipped solar sail, accounting for the effect of changing the solar sail’s CM. This provides an accurate prediction model that can be used within an MPC policy. The motion of the AMT is then discretized with a first-order-hold (FOH) allowing for piece-wise linear motion of the AMT to be accurately captured in the prediction model. The RCDs operate in an on-off fashion, which traditionally would manifest as integer variables within an MPC optimization scheme. This type of optimization problem is highly non-convex and is usually solved using dynamic programming, mixed-integer programming, or branch-and-bound approach [30, 31, 32, 33], all of which demand significant computational memory and time. To avoid this impractical approach, we propose solving an initial MPC problem where the RCD torque is allowed to vary continuously between its positive and negative actuation magnitude, then quantizing the result into a single on-off pulse. This initial problem can be solved as a quadratic problem (QP) in an efficient fashion using a linear approximation of the system’s dynamics. A pulse-width-modulation (PWM) quantization inspired by the thruster momentum management actuation in [34, 35, 36, 26, 33] is then applied to turn the optimal continuous RCD input solution from MPC into a single discrete pulse with varying pulse length (duration) within a single timestep. The result is an MPC policy that can be computed efficiently and reliable, while ensuring a physically-realizable quantized RCD input.
@@
A second version of the proposed MPC-based momentum management strategy is presented in this work to address the fact that performing PWM quantization on the MPC solution may result sub-optimal performance. This is addressed through a novel iterative backwards-in-time MPC strategy that fixes the last RCD input in the prediction horizon as a quantized value and re-optimizes the remainder of the control inputs. This process is repeated by fixing the next-to-last RCD input as a quantized value, and continuing this process until eventually the first RCD input is quantized. This iterative approach removes the complexity of of a branch-and-bound approach to integer programming, while including the quantization knowledge in the MPC prediction model to obtain improved performance.
@@
The contributions of this paper are highlighted as 1) the dynamic modeling of a solar sail equipped with AMT accounting for the effects of a moving CM; 2) two solar sail momentum management strategies using MPC that meet practical operational needs through the explicit inclusion of actuator magnitude, actuator rate limits, and quantization of the RCD inputs, which is shown to significantly reduce the actuation required compared to the state-of-the-art method in [11, 16]; and 3) a novel backwards-in-time iterative solution approach to the second MPC strategy that incorporates knowledge of the quantized RCD input into the MPC optimization problem in a computationally-efficient manner by solving a sequence of convex optimization problems. The practical challenges associated with the momentum management actuators and a detailed derivation of the dynamic model are presented in Section 2. In Section 3, an MPC-based momentum management controller is formulated to meet the operational requirements of the solar sail and two different methods are proposed to address the quantized RCD inputs. Numerical simulations and comparisons between the proposed methods using practical solar sail parameters are presented in Section 4.
@@ 2 System Dynamics
Before deriving the detailed solar sail dynamics, Section 2.1 provides a brief overview of the momentum management actuators utilized in NASA’s Solar Cruiser: the AMT and RCDs. A derivation of the solar sail’s attitude dynamics, accounting for the effects of AMT translation on the spacecraft’s CM and moment of inertia, is presented in Sections 2.2 and 2.3. A linearization procedure for the derived nonlinear model is detailed in Section 3.1, serving as the prediction model for MPC. The nonlinear model itself is subsequently used in the numerical simulations of Section 4. Unlike conventional rigid-body spacecraft dynamics, the solar sail introduces unique challenges. The AMT’s ability to translate a significant portion of the spacecraft’s mass creates dynamics that are directly dependent on the AMT’s position. This dependency is highlighted throughout the modeling and control synthesis processes in Sections 2.2 and 3.
@@ 2.1 Momentum Management Actuators
The focus of this paper is to design a momentum management strategy that efficiently unloads the angular momentum accumulated in the RWs. To achieve this, AMT and RCDs are employed as the primary momentum management actuators on NASA’s Solar Cruiser [11].
An active mass translator (AMT), depicted in Fig. 1, moves a portion of the spacecraft bus mass within a plane parallel to the sail surface. By adjusting the AMT’s position, the relative positions of the CM and the CP are modified, creating a moment arm when SRP acts on the sail surface. This generates torques about the pitch and yaw axes (respectively denoted as the $\underrightarrow{b}^{2}$ and $\underrightarrow{b}^{1}$ axes in Section 2.2). Solar Cruiser’s AMT translates approximately 5050% of the spacecraft’s mass, enabling a substantial CM displacement within the AMT travel range of $\pm 0.29$ m [4]. This torque can mitigate RW angular momentum buildup and counteract disturbances by moving the AMT to the appropriate position. However, the CM displacement caused by AMT motion also alters the spacecraft’s moment of inertia and overall dynamics. The derivation of a dynamic model accounting for these effects is detailed in Section 2.2. In this paper, the AMT’s control input is modeled to reflect realistic motion constraints. Instead of a zeroth-order hold (ZOH), which assumes abrupt positional changes, a FOH is used to simulate continuous translation between momentum management time steps. This representation also facilitates the inclusion of AMT translation rate limits as constraints in the MPC framework. Details of the discretization and linear interpolation are presented in Section 3.2.
@@
Reflectivity control devices (RCDs) serve as the key momentum management actuators for the roll axis (denoted as the $\underrightarrow{b}^{3}$ axis in Section 2.2). Each RCD operates by modulating its reflectivity using small electrical power inputs, altering the distribution of SRP on the sail and generating torques for roll-axis control. Pairs of RCDs are strategically positioned near the outer edges of each sail boom, as illustrated in Fig. 2. These pairs are arranged in an inclined “tent-like” configuration with opposing angles, creating a differential reflectivity that produces roll-axis torque. This torque, denoted as ${\bm{\tau}}_{b}^{\text{RCD}}$, is treated as a pure roll-axis control input in the dynamic model described in Section 2.2. Disturbances arising from sail surface imperfection, roughness, or deformation are modeled separately as disturbance torque ${\bm{\tau}}_{b}^{\text{dist}}$ in Section 2.2. The unique actuation mechanism of RCDs requires careful modeling due to their discrete, electrically powered operation. They function in on-off actuation pulses, represented by integer values $\alpha_{\text{on-off}}\in\{-1,0,1\}$ corresponding to negative, zero, and positive directions of a fixed magnitude RCDs torque ${\bm{\tau}}_{b,\text{on}}^{\text{RCD}}$, i.e., ${\bm{\tau}}_{b}^{\text{RCD}}=\alpha_{\text{on-off}}\cdot{\bm{\tau}}_{b,\text{on}}^{\text{RCD}}$. This discrete nature introduces a non-convex integer constraint, presenting computational challenges for efficient control optimization. Approaches to address these constraints, including PWM-quantization and iterative backtracking methods, are discussed in Sections 3.3 and 3.5.
@@ 2.2 Nonlinear System Dynamics
Let $\mathcal{F}_{a}$, defined by basis vectors $\underrightarrow{a}^{1}$, $\underrightarrow{a}^{2}$, $\underrightarrow{a}^{3}$, be the inertial reference frame. A solar sail body-fixed frame $\mathcal{F}_{b}$ is defined as $\underrightarrow{b}^{3}$ pointing through the roll axis of the sail, and $\underrightarrow{b}^{2}$, $\underrightarrow{b}^{1}$ denote the pitch and yaw axes, respectively. A 3-2-1 Euler angle sequence is used to describe the rotation between $\mathcal{F}_{a}$ and $\mathcal{F}_{b}$, so that $\mbf{C}_{ba}=\mbf{C}_{1}(\theta_{1})\mbf{C}_{2}(\theta_{2})\mbf{C}_{3}(\theta_{3})$ is the direction cosine matrix (DCM) describing the orientation of $\mathcal{F}_{b}$ relative to $\mathcal{F}_{a}$, and ${\bm{\omega}}^{ba}_{b}=\Big{[}\omega^{ba}_{b1}\,\,\,\omega^{ba}_{b2}\,\,\,\omega^{ba}_{b3}\Big{]}^{\top}=\mbf{S}({\bm{\theta}})\dot{{\bm{\theta}}}$ is the angular velocity of $\mathcal{F}_{b}$ relative to $\mathcal{F}_{a}$ resolved in $\mathcal{F}_{b}$, where ${\bm{\theta}}=\Big{[}\theta_{1}\,\,\,\theta_{2}\,\,\,\theta_{3}\Big{]}^{\top}$ is the set of yaw, pitch, and roll Euler angles, and $\mbf{S}({\bm{\theta}})$ is the mapping matrix between angular rates $\dot{{\bm{\theta}}}$, given by
@@
Consider a rigid body spacecraft $\mathcal{B}$, as shown in Fig. 3, consisting of a square planar sail membrane $\mathcal{S}$ with mass $m_{s}$, and a rectangular cuboid bus $\mathcal{P}$ with mass $m_{p}$. For simplicity, it is assumed that both $\mathcal{S}$ and $\mathcal{P}$ have their CM colocated with their geometric center, denoted by points $s$ and $p$. The CM of the entire spacecraft $\mathcal{B}$ is denoted by point $c$.
@@
The bus $\mathcal{P}$ houses the attitude control system with three RWs located at point $p$, which is coincident with the CM of the bus. Each of the RWs spins about one of the three nominal axes $\underrightarrow{b}^{1}$, $\underrightarrow{b}^{2}$, $\underrightarrow{b}^{3}$ of the spacecraft body frame $\mathcal{F}_{b}$. These assumptions can be relaxed in principle, but are chosen to simplify portions of the system dynamics in this work. Point $p$ relative to point $s$ resolved in frame $\mathcal{F}_{b}$ is denoted as $\mbf{r}^{ps}_{b}$. This position vector is determined by the AMT’s translation of the bus $\mathcal{P}$ relative to the sail $\mathcal{S}$ within the ${\underrightarrow{{b}}}^{1}-{\underrightarrow{{b}}}^{2}$ plane, such that $\mbf{r}^{ps}_{b}(t)=\Big{[}r^{\text{AMT}}_{b1}(t)\,\,\,r^{\text{AMT}}_{b2}(t)\,\,\,{r}^{ps}_{b3}\Big{]}^{\top}$, where the distance ${r}^{ps}_{b3}$ is constant.
@@
$\mbf{J}_{b}^{\mathcal{P}c}(t)=\mbf{J}^{\mathcal{P}p}_{b}-m_{p}\mbf{r}_{b}^{{pc}^{\times}}(t)\mbf{r}_{b}^{{pc}^{\times}}(t),$
@@
[…]
$\mbf{u}_{b}^{\times}=\begin{bmatrix}u_{b1}\\ u_{b2}\\ u_{b3}\end{bmatrix}^{\times}=\begin{bmatrix}0&-u_{b3}&u_{b2}\\ u_{b3}&0&-u_{b1}\\ -u_{b2}&u_{b1}&0\end{bmatrix}$.
The moment of inertia of the whole spacecraft $\mathcal{B}$ relative to the collective CM of the spacecraft (point $c$ in Fig. 3) is
$\mbf{J}_{b}^{\mathcal{B}c}(t)=\mbf{J}_{b}^{\mathcal{P}c}(t)+\mbf{J}_{b}^{\mathcal{S}c}(t)=\mbf{J}^{\mathcal{P}p}_{b}+\mbf{J}^{\mathcal{S}s}_{b}-\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}\mbf{r}_{b}^{{ps}^{\times}}(t)\mbf{r}_{b}^{{ps}^{\times}}(t).$
@@
$\mbf{h}_{b}^{\mathcal{S}c/a}(t)$ $=\mbf{J}^{\mathcal{S}c}_{b}(t){\bm{\omega}}^{ba}_{b}(t)+m_{s}\mbf{r}^{{sc}^{\times}}_{b}(t)\dot{\mbf{r}}^{sc}_{b}(t),$
$\mbf{h}_{b}^{\mathcal{G}c/a}(t)$ $=\mbf{J}^{\mathcal{G}c}_{b}(t){\bm{\omega}}^{ba}_{b}(t)+m_{g}\mbf{r}^{{pc}^{\times}}_{b}(t)\dot{\mbf{r}}^{pc}_{b}(t).$
@@
$\mbf{h}_{b}^{\mathcal{W}_{i}c/a}(t)=\mbf{J}^{\mathcal{W}_{i}p}_{b}\dot{{\bm{\gamma}}}^{\mathcal{W}_{i}}(t)+m_{w}\mbf{r}^{pc^{\times}}_{b}(t)\dot{\mbf{r}}^{pc}_{b}(t)+\mbf{J}^{\mathcal{W}_{i}c}_{b}(t)\dot{{\bm{\gamma}}}^{\mathcal{W}_{i}}(t),\hskip 11.49994pti=1,2,3,$

### v3

# Solar Sail Momentum Management With Mass Translation and Reflectivity Devices Using Predictive Control

@@ Abstract
keywords:
Solar Sails , Momentum Management , Model Predictive Control , Active Mass Translator , Reflectivity Control Devices
journal: Acta Astronautica
@@ 1 Introduction
Space sailing is a field of growing interest with diverse applications in space science and engineering [1]. Considering the growing number of space debris [2, 3], drag sails have been proposed for the deorbiting end-of-life satellites [4, 5, 6]. On the other hand, the concept of a solar sail introduces the opportunity for fuel-efficient, and potentially even fuel-free, inter-planetary travel and deep-space exploration [7, 8, 9]. The momentum transferred to a solar sail by solar radiation pressure (SRP) provides a promising source of propulsive thrust that can be controlled by adjusting the attitude of the solar sail, and thus the direction and magnitude of thrust. A larger area-to-mass ratio increases a solar sail’s efficiency, which motivates the desire of developing a large-scale solar sail system, such as NASA’s Solar Cruiser, which involves a sail area greater than $1600$ m2 [10, 11, 12, 13]. However, larger-scale lightweight solar sails are inherently flexible, resulting in undesirable structural deformations that shift the solar sail’s center of pressure (CP) and create disturbance torques that need to be rejected by an attitude control system [14, 15, 16, 17, 18, 19]. Reaction wheels (RWs) are typically used for this purpose, but eventually suffer from an accumulation of stored angular momentum, which causes a loss of attitude control authority when the RWs saturate, necessitating momentum management. Standard momentum management strategies for spacecraft, such as the use of thrusters or magnetic actuation, are not well-suited for solar sails, since they either require onboard fuel or the presence of an external magnetic field. To truly take advantage of the potential of solar sails for fuel-free deep-space exploration, more creative momentum management solutions are required.
@@
To solve this challenge, Solar Cruiser uses an active mass translator (AMT) and reflectivity control devices (RCDs) as actuators to respectively generate pitch/yaw and roll torques for momentum management [20, 21]. The AMT serves as a mechanism between two portions of the spacecraft bus, allowing them to translate relative to each other in a plane. This results in a controllable shift of the solar sail’s center of mass (CM), allowing for the generation of SRP-induced torques or the “trimming” out of disturbance torques in the pitch and yaw axes. Unlike rotative tip-devices [22, 23, 24, 25], the RCDs are stationary thin film membranes located towards the extremities of the solar sail membrane and are canted at an angle to the plane of the sail membrane. They generate roll torques about the axis normal to the sail by adjusting their reflectivity when a voltage is applied to them. A challenge in the operation of RCDs is that they operate in an on-off fashion, either generating no torque or a constant magnitude torque in the positive or negative roll direction. Solar Cruiser’s attitude determination and control system includes control subsystems for momentum management that command the AMT and RCDs [26]. The current design of this momentum management system involves decoupled proportional-integral-derivative (PID) controllers to actuate the two axes of the AMT based on the momentum stored in the pitch/yaw axis RWs, as well as an on-off actuation strategy for the RCDs [20, 26]. Although this simple control strategy was shown to successfully perform momentum management, it neglects the dynamic coupling between the AMT and RCD actuators, which can result in undesirable performance bordering on closed-loop instability [20]. Additionally, the on-off nature of the RCDs is handled in the approach from [20, 26] using activation/deactivation thresholds, which creates additional nonlinearities in the feedback system, affecting closed-loop performance and stability properties. There is a pressing need for a momentum management control policy that is capable of explicitly accounting for the nonlinear, coupled dynamics involved in the AMT and RCD actuators, as well as the practical constraints associated with their operation.
@@
In this work, MPC is proposed for solar sail momentum management as a means to account for the environmental disturbance torques acting on the solar sail, as well as the coupled nature of the actuators and the nonlinearities of the system’s dynamics. Given the limited flight hardware and computational resources onboard, a simple dynamic model and convex formulation are necessary for real-time implementation. The unique actuation properties of the AMT and RCDs make this MPC formulation challenging, resulting in an optimization problem that is naturally non-convex and unsuitable for implementation on flight hardware. To remedy this issue, we first derive a dynamic model of an AMT equipped solar sail, accounting for the effect of changing the solar sail’s CM. This provides an accurate prediction model that can be used within an MPC policy. The motion of the AMT is then discretized with a first-order-hold (FOH) allowing for piece-wise linear motion of the AMT to be accurately captured in the prediction model. The RCDs operate in an on-off fashion, which traditionally would manifest as integer variables within an MPC optimization scheme. This type of optimization problem is highly non-convex and is usually solved using dynamic programming, mixed-integer programming, or branch-and-bound approach [40, 41, 42, 43], all of which demand significant computational memory and time. To avoid this impractical approach, we propose solving an initial MPC problem where the RCD torque is allowed to vary continuously between its positive and negative actuation magnitude, then quantizing the result into a single on-off pulse. This initial problem can be solved as a quadratic problem (QP) in an efficient fashion using a linear approximation of the system’s dynamics. A pulse-width-modulation (PWM) quantization inspired by the thruster momentum management actuation in [44, 45, 46, 36, 43] is then applied to turn the optimal continuous RCD input solution from MPC into a single discrete pulse with varying pulse length (duration) within a single timestep. The result is an MPC policy that can be computed efficiently and reliable, while ensuring a physically-realizable quantized RCD input.
@@
A second version of the proposed MPC-based momentum management strategy is presented in this work to address the fact that performing PWM quantization on the MPC solution may result sub-optimal performance. This is addressed through a novel iterative backwards-in-time MPC strategy that fixes the last RCD input in the prediction horizon as a quantized value and re-optimizes the remainder of the control inputs. This process is repeated by fixing the next-to-last RCD input as a quantized value, and continuing this process until eventually the first RCD input is quantized. This iterative approach removes the complexity of a branch-and-bound approach to integer programming, while including the quantization knowledge in the MPC prediction model to obtain improved performance.
@@
The contributions of this paper are highlighted as 1) the dynamic modeling of a solar sail equipped with AMT accounting for the effects of a moving CM; 2) two solar sail momentum management strategies using MPC that meet practical operational needs through the explicit inclusion of actuator magnitude, actuator rate limits, and quantization of the RCD inputs, which is shown to significantly reduce the actuation required compared to the state-of-the-art method in [20, 26]; and 3) a novel backwards-in-time iterative solution approach to the second MPC strategy that incorporates knowledge of the quantized RCD input into the MPC optimization problem in a computationally-efficient manner by solving a sequence of convex optimization problems. Together, these contributions represent an advance in state-of-the-art solar sail AMT/RCD momentum management, which up until now has struggled to meet operational requirements due to the use of decoupled PID control loops with on-off thresholding. The practical challenges associated with the momentum management actuators and a detailed derivation of the dynamic model are presented in Section 3. In Section 4, an MPC-based momentum management controller is formulated to meet the operational requirements of the solar sail and two different methods are proposed to address the quantized RCD inputs. Numerical simulations and comparisons between the proposed methods using practical solar sail parameters are presented in Section 5.
2 Notation
In this paper, reference frame $\mathcal{F}_{a}$ is defined by three orthonormal, dextral basis vectors $\underrightarrow{a}^{1}$, $\underrightarrow{a}^{2}$, $\underrightarrow{a}^{3}$. The position vector $\underrightarrow{r}^{ij}$ describes the position of point $i$ relative to point $j$ and its components when resolved in reference frame $\mathcal{F}_{a}$ are denoted by $\mbf{r}^{ij}_{a}=\begin{bmatrix}r^{ij}_{a1}&r^{ij}_{a2}&r^{ij}_{a3}\end{bmatrix}^{\top}$. The cross product $\underrightarrow{w}=\underrightarrow{u}\times\underrightarrow{v}$ resolved in $\mathcal{F}_{a}$ is expressed as $\mbf{u}^{\times}_{a}\mbf{v}_{a}$, where
$\mbf{u}^{\times}_{a}=\begin{bmatrix}u_{a1}\\ u_{a2}\\ u_{a3}\end{bmatrix}^{\times}=\begin{bmatrix}0&-u_{a3}&u_{a2}\\ u_{a3}&0&-u_{a1}\\ -u_{a2}&u_{a1}&0\end{bmatrix}$ (1)
is the skew-symmetric cross-product matrix of $\mbf{u}_{a}$. The direction cosine matrix (DCM) $\mbf{C}_{ba}$ describes the orientation of reference frame $\mathcal{F}_{b}$ relative to $\mathcal{F}_{a}$ and can be used to convert a vector resolved in one frame to another (e.g., $\mbf{u}_{b}=\mbf{C}_{ba}\mbf{u}_{a}$).
@@ 3 System Dynamics
Before deriving the detailed solar sail dynamics, Section 3.1 provides an overview of the momentum management actuators utilized in NASA’s Solar Cruiser: the AMT and RCDs. A derivation of the solar sail’s attitude dynamics, accounting for the effects of AMT translation on the spacecraft’s CM and moment of inertia, is presented in Sections 3.2 and 3.3. A linearization procedure for the derived nonlinear model is detailed in Section 4.1, serving as the prediction model for MPC. The nonlinear model itself is subsequently used in the numerical simulations of Section 5. Unlike conventional rigid-body spacecraft dynamics, the solar sail introduces unique challenges. The AMT’s ability to translate a significant portion of the spacecraft’s mass creates dynamics that are directly dependent on the AMT’s position. This dependency is highlighted throughout the modeling and control synthesis processes in Sections 3.2 and 4.
@@ 3.1 Momentum Management Actuators
The focus of this paper is to design a momentum management strategy that efficiently unloads the angular momentum accumulated in the RWs. To achieve this, AMT and RCDs are employed as the primary momentum management actuators on NASA’s Solar Cruiser [20], which are scaled to perform control authority >1.5>1.5 times of the disturbance torque [13].
An active mass translator (AMT), depicted in Figure 1, moves a portion of the spacecraft bus mass within a plane parallel to the sail surface. By adjusting the AMT’s position, the relative positions of the CM and the CP are modified, creating a moment arm when SRP acts on the sail surface. This generates torques about the pitch and yaw axes (respectively denoted as the $\underrightarrow{b}^{2}$ and $\underrightarrow{b}^{1}$ axes in Section 3.2). Solar Cruiser’s AMT translates approximately 5050% of the spacecraft’s mass, enabling a substantial CM displacement within the AMT travel range of $\pm 0.29$ m [10]. This torque can mitigate RW angular momentum buildup and counteract disturbances by moving the AMT to the appropriate position. However, the CM displacement caused by AMT motion also alters the spacecraft’s moment of inertia and overall dynamics. The derivation of a dynamic model accounting for these effects is detailed in Section 3.2. In this paper, the AMT’s control input is modeled to reflect realistic motion constraints. Instead of a zeroth-order hold (ZOH), which assumes abrupt positional changes, a FOH is used to simulate continuous translation between momentum management time steps. This representation also facilitates the inclusion of AMT translation rate limits as constraints in the MPC framework. Details of the discretization and linear interpolation are presented in Section 4.2.
@@
Reflectivity control devices (RCDs) serve as the key momentum management actuators for the roll axis (denoted as the $\underrightarrow{b}^{3}$ axis in Section 3.2). Each RCD operates by modulating its reflectivity using small electrical power inputs, altering the distribution of SRP on the sail and generating torques for roll-axis control. A total of eight RCDs are strategically positioned in pairs near the outer edges of each sail boom, as illustrated in Figure 2. These pairs are arranged in an inclined “tent-like” configuration with opposing angles. Turning on four of the RCDs with the same tent angle induces a roll-axis torque due to an imbalance in the differential forces in-plane to the sail. The differential out-of-plane forces induced by each of the RCDs are canceled out due to symmetry. This torque, denoted as ${\bm{\tau}}_{b}^{\text{RCD}}$, is treated as a pure roll-axis control input in the dynamic model described in Section 3.2. Disturbances arising from sail surface imperfection, roughness, or deformation are modeled separately as disturbance torque ${\bm{\tau}}_{b}^{\text{dist}}$ in Section 3.2. The unique actuation mechanism of RCDs requires careful modeling due to their discrete, electrically powered operation. They function in on-off actuation pulses, represented by integer values $\alpha_{\text{on-off}}\in\{-1,0,1\}$ corresponding to negative, zero, and positive directions of a fixed magnitude RCDs torque ${\bm{\tau}}_{b,\text{on}}^{\text{RCD}}$, i.e., ${\bm{\tau}}_{b}^{\text{RCD}}=\alpha_{\text{on-off}}\cdot{\bm{\tau}}_{b,\text{on}}^{\text{RCD}}$. This discrete nature introduces a non-convex integer constraint, presenting computational challenges for efficient control optimization. Approaches to address these constraints, including PWM-quantization and iterative backtracking methods, are discussed in Sections 4.3 and 4.5.
@@ 3.2 Nonlinear System Dynamics
Let $\mathcal{F}_{a}$, defined by basis vectors $\underrightarrow{a}^{1}$, $\underrightarrow{a}^{2}$, $\underrightarrow{a}^{3}$, be the inertial reference frame. A solar sail body-fixed frame $\mathcal{F}_{b}$ is defined as $\underrightarrow{b}^{3}$ pointing through the roll axis of the sail, and $\underrightarrow{b}^{2}$, $\underrightarrow{b}^{1}$ denote the pitch and yaw axes, respectively. A 3-2-1 Euler angle sequence is used to describe the rotation between $\mathcal{F}_{a}$ and $\mathcal{F}_{b}$, so that $\mbf{C}_{ba}=\mbf{C}_{1}(\theta_{1})\mbf{C}_{2}(\theta_{2})\mbf{C}_{3}(\theta_{3})$ is the DCM describing the orientation of $\mathcal{F}_{b}$ relative to $\mathcal{F}_{a}$, and ${\bm{\omega}}^{ba}_{b}=\Big[\omega^{ba}_{b1}\,\,\,\omega^{ba}_{b2}\,\,\,\omega^{ba}_{b3}\Big]^{\top}=\mbf{S}({\bm{\theta}})\dot{{\bm{\theta}}}$ is the angular velocity of $\mathcal{F}_{b}$ relative to $\mathcal{F}_{a}$ resolved in $\mathcal{F}_{b}$, where ${\bm{\theta}}=\Big[\theta_{1}\,\,\,\theta_{2}\,\,\,\theta_{3}\Big]^{\top}$ is the set of yaw, pitch, and roll Euler angles, and $\mbf{S}({\bm{\theta}})$ is the mapping matrix between angular rates $\dot{{\bm{\theta}}}$, given by
@@
Consider a rigid body spacecraft $\mathcal{B}$, as shown in Figure 3, consisting of a square planar sail membrane $\mathcal{S}$ with mass $m_{s}$, and a rectangular cuboid bus $\mathcal{P}$ with mass $m_{p}$. The dimension of bus $\mathcal{P}$ is given by $\ell_{1}$, $\ell_{2}$, $\ell_{3}$ in each edge. The square sail $\mathcal{S}$ has length $L$ in each edge. For simplicity, it is assumed that both $\mathcal{S}$ and $\mathcal{P}$ have their CM colocated with their geometric center, denoted by points $s$ and $p$. The CM of the entire spacecraft $\mathcal{B}$ is denoted by point $c$.
@@
The bus $\mathcal{P}$ houses the attitude control system with three RWs (denoted by $\mathcal{W}_{1}$, $\mathcal{W}_{2}$, $\mathcal{W}_{3}$) located at point $p$, which is coincident with the CM of the bus. Each of the RWs spins about one of the three nominal axes $\underrightarrow{b}^{1}$, $\underrightarrow{b}^{2}$, $\underrightarrow{b}^{3}$ of the spacecraft body frame $\mathcal{F}_{b}$. These assumptions can be relaxed in principle, but are chosen to simplify portions of the system dynamics in this work. Point $p$ relative to point $s$ resolved in frame $\mathcal{F}_{b}$ is denoted as $\mbf{r}^{ps}_{b}$. This position vector is determined by the AMT’s translation of the bus $\mathcal{P}$ relative to the sail $\mathcal{S}$ within the ${\underrightarrow{{b}}}^{1}-{\underrightarrow{{b}}}^{2}$ plane, such that $\mbf{r}^{ps}_{b}(t)=\Big[r^{\text{AMT}}_{b1}(t)\,\,\,r^{\text{AMT}}_{b2}(t)\,\,\,{r}^{ps}_{b3}\Big]^{\top}$, where the distance ${r}^{ps}_{b3}$ is constant.
@@
$\mbf{J}_{b}^{\mathcal{P}c}(t)=\mbf{J}^{\mathcal{P}p}_{b}-m_{p}\mbf{r}_{b}^{{pc}^{\times}}(t)\mbf{r}_{b}^{{pc}^{\times}}(t),$ (2)
@@
where $\mbf{r}_{b}^{pc}(t)=-\frac{m_{p}}{m_{p}+m_{s}}\mbf{r}^{ps}_{b}(t)$ and $\mbf{r}_{b}^{sc}(t)=\frac{m_{s}}{m_{p}+m_{s}}\mbf{r}^{ps}_{b}(t)$ depend on the mass ratio between bus $\mathcal{P}$ and sail $\mathcal{S}$, as well as the position of the AMT. The moment of inertia of the whole spacecraft $\mathcal{B}$ relative to the collective CM of the spacecraft (point $c$ in Figure 3) is
$\mbf{J}_{b}^{\mathcal{B}c}(t)=\mbf{J}_{b}^{\mathcal{P}c}(t)+\mbf{J}_{b}^{\mathcal{S}c}(t)=\mbf{J}^{\mathcal{P}p}_{b}+\mbf{J}^{\mathcal{S}s}_{b}-\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}\mbf{r}_{b}^{{ps}^{\times}}(t)\mbf{r}_{b}^{{ps}^{\times}}(t).$ (4)
@@
$\mbf{h}_{b}^{\mathcal{S}c/a}(t)$ $=\mbf{J}^{\mathcal{S}c}_{b}(t){\bm{\omega}}^{ba}_{b}(t)+m_{s}\mbf{r}^{{sc}^{\times}}_{b}(t)\dot{\mbf{r}}^{sc}_{b}(t),$ (5)
$\mbf{h}_{b}^{\mathcal{G}c/a}(t)$ $=\mbf{J}^{\mathcal{G}c}_{b}(t){\bm{\omega}}^{ba}_{b}(t)+m_{g}\mbf{r}^{{pc}^{\times}}_{b}(t)\dot{\mbf{r}}^{pc}_{b}(t).$ (6)
@@
$\mbf{h}_{b}^{\mathcal{W}_{i}c/a}(t)=\mbf{J}^{\mathcal{W}_{i}p}_{b}\dot{{\bm{\gamma}}}^{\mathcal{W}_{i}}(t)+m_{w}\mbf{r}^{pc^{\times}}_{b}(t)\dot{\mbf{r}}^{pc}_{b}(t)+\mbf{J}^{\mathcal{W}_{i}c}_{b}(t)\dot{{\bm{\gamma}}}^{\mathcal{W}_{i}}(t),\hskip 11.49994pti=1,2,3,$ (7)